OpenStudy (ashley1nonly):

take the anti-derivative of sin(1/7x) dx

OpenStudy (ashley1nonly):

i know that we have to use the chain rule

OpenStudy (danjs):

\[\int\limits \sin(\frac{ 1 }{ 7 }x)dx\]

OpenStudy (misty1212):

backwards

OpenStudy (ashley1nonly):

since it sin it would have to go to -cos

OpenStudy (misty1212):

you want to say \(-\cos(\frac{x}{7})\) but that is wrong because you would get \[\frac{1}{7}\sin(\frac{x}{7})\] by the chain rule no worries, multiply by 7 to fix is

OpenStudy (ashley1nonly):

that what i dont understant. where is the extra 7 coming from

OpenStudy (danjs):

let u = 1/7 x du = 1/7 dx so dx = 7 du \[\int\limits \sin(u) *7 *du\] integrate, then sub back in what x is

OpenStudy (ashley1nonly):

this is substitution right, i didnt think you could use it for this one

OpenStudy (danjs):

yeah, u - sub, so you get \[7*\int\limits \sin(u) du = 7*-\cos(u) \] u = 1/7 x

OpenStudy (danjs):

-7 cos(1/7 x)

OpenStudy (ashley1nonly):

then you pulled the 7 out because it was a constant right

OpenStudy (danjs):

right, you let u = 1/7x and the derivative is du = 1/7 dx solve that for dx, dx = 7 du pull the constant outside the integral

OpenStudy (misty1212):

the extra 7 from the chain rule if you take the derivative you have to divide by 7, so when you take the anti derivative you have to multiply by 7

OpenStudy (ashley1nonly):

-7cos(x/7) + c

OpenStudy (ashley1nonly):

okay why did i leave the du alone in the equation for example 7/ sin(u) du

OpenStudy (ashley1nonly):

wouldnt i have to take it back to dx or no

OpenStudy (danjs):

you make a substitution, so your integral is in U-space , you complete the integral in U-space, then substitute back in at the end

OpenStudy (danjs):

you defined u = 1/7 x

OpenStudy (ashley1nonly):

Okay i get that, thank you so much

OpenStudy (danjs):

welcome